OpenStudy (anonymous):

Find the vertex form of x^2+6x+5=y

OpenStudy (anonymous):

@iambatman @amistre64 @jdoe0001 @kohai

OpenStudy (anonymous):

I got (x+30^2=(y+4) this is the first day we've learned about the topic so i want to make sure i'm doing this right

OpenStudy (anonymous):

sorry, (x+3)^2

OpenStudy (imstuck):

This is a parabola. The way you find the vertex form of this is to complete the square on the x terms. Do you know how to do that? Let me check your answer. Give me a sec, ok?

OpenStudy (imstuck):

the answer is either this:\[(x+3)^{2}=y+4\]or\[(x+3)^{2}-4=y\]Either way, you are correct. Good job!

OpenStudy (anonymous):

if i have it in the form (x-h)^2=4p(y-k) what is p if there isnt a number there? 1/4?

OpenStudy (jdoe0001):

\(\bf x^2+6x+5=y\implies (x+3)^2-4=y\implies (x+3)^2=1(y+4) \\ \quad \\ \begin{cases} 4p=1\\ p=\frac{1}{4} \end{cases}\implies (x-({\color{brown}{ -3}})^2=4\cdot {\color{blue}{ \frac{1}{4}}}(y-({\color{brown}{ -4}}))\)